OpenStudy (anonymous):

Assuming that v is in the interval (0,pi/2) and sin v=1/4 evaluate cos(2v) Enter an exact answer=?

OpenStudy (anonymous):

Since v is in the first quadrant, cos v must be positive. We know that \[\sin ^{2}v + \cos^2v = 1\]. Therefore cos v = \[\sqrt{15}/4\]. Since cos2v = (cosv)^2 - (sinv)^2. We can find the answer.

OpenStudy (anonymous):

thanks